OpenStudy (kayders1997):

Please help will fan and medal, find the limit using l'hospital's rule (xe^3x-x)/(1-cos2x)

OpenStudy (freckles):

x approaches 0?

OpenStudy (kayders1997):

my bad im sorry the limit does approach 0

OpenStudy (freckles):

\[\lim_{x \rightarrow 0} \frac{xe^{3x}-x}{1-\cos(2x)}\]

OpenStudy (kayders1997):

yes

OpenStudy (freckles):

so we do have 0/0 when x approaches 0 so that does mean we can apply l'hospital

OpenStudy (kayders1997):

yes

OpenStudy (freckles):

have you tried differentiating top and bottom w.r.t. x ?

OpenStudy (kayders1997):

I can do the bottom

OpenStudy (kayders1997):

and I know -x=-1 for the top

OpenStudy (freckles):

\[(xe^{3x}-x)' \\ =(xe^{3x})'-(x)' \text{ by difference rule } \\ =(x)'e^{3x}+x(e^{3x})'-(x)' \text{ by product rule } \\ =...\]

OpenStudy (freckles):

so you are left differentiating x and e^(3x)

OpenStudy (kayders1997):

yes so it would be? e^3x+xe^3x-1?

OpenStudy (freckles):

almost you have differentiated e^(3x) just a bit incorrectly

OpenStudy (kayders1997):

Oh wait would it be 3e^3x?

OpenStudy (freckles):

yep

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{e^{3x}+3xe^{3x}-1}{2 \sin(2x)}\]

OpenStudy (kayders1997):

okay I'm good to this step

OpenStudy (freckles):

nothing is obvious to me here but when I do plug in 0 again I do get 0/0 so I would try applying l'hospital again

OpenStudy (kayders1997):

omg!!! this is going to be really messy

OpenStudy (freckles):

but...

OpenStudy (freckles):

notice the bottom will be 4 cos(2x) and 4 cos(2x) evaluated at x=0 is not 0 :) so we are almost done

OpenStudy (kayders1997):

yay :)

OpenStudy (freckles):

try diff the numerator

OpenStudy (kayders1997):

okay

OpenStudy (kayders1997):

well I know the first part is 3e^3x

OpenStudy (freckles):

for the second term you will have to apply product rule

OpenStudy (kayders1997):

Okay

OpenStudy (freckles):

\[(3xe^{3x})' \\ (3x)' e^{3x}+3x(e^{3x})' \text{ by product rule }\]

OpenStudy (kayders1997):

so? 3(x)'(e^3x)+(e^3x)'(x)

OpenStudy (freckles):

almost you missed the 3 in the second term you just wrote

OpenStudy (freckles):

like a factor of 3 that is

OpenStudy (kayders1997):

oops

OpenStudy (kayders1997):

so than you will get 3e^3x+9xe^3x?

OpenStudy (freckles):

one sec let me check

OpenStudy (freckles):

\[(e^{3x}+3x e^{3x}-1)' \\ =(e^{3x})'+(3xe^{3x})'-(1)' \\ =3e^{3x}+3e^{3x}+3x \cdot 3e^{3x}-0 \\ =6e^{3x}+9xe^{3x}\]

OpenStudy (freckles):

oh you mean just for the second term you are right

OpenStudy (freckles):

and then we add the derivative from that first term and we are good to go

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{6e^{3x}+9xe^{3x}}{4 \cos(2x)}\]

OpenStudy (freckles):

now you are sorta done with the calculus part

OpenStudy (freckles):

just need to plug in 0

OpenStudy (kayders1997):

okay

OpenStudy (kayders1997):

e^0 is 1 right?

OpenStudy (freckles):

yep

OpenStudy (kayders1997):

I got 3/2

OpenStudy (freckles):

:)

OpenStudy (freckles):

looks great you get 6/4 which reduces to 3/2 great job

OpenStudy (kayders1997):

thank you :) can I ask you about one more but I'll put it in a new tab, it was one we had to do in class but I didn't get it right

OpenStudy (freckles):

sure

OpenStudy (kayders1997):

Thank you :)